Circular No. 7143

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999br IN NGC 4900 P. Garnavich, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999br was obtained by M. Calkins with the 1.5-m Tillinghast telescope on Apr. 15.3 UT. Broad Balmer emission lines on a very blue continuum are seen, indicating that this is a type-II event discovered at an early stage. Hα shows only a weak absorption trough, but the later Balmer series have clear P-Cyg profiles. He I 587.5-nm has an emission flux comparable to that of Hβ. Narrow host-galaxy emission lines give a recession velocity of 965 km/s.

A. V. Filippenko, University of California at Berkeley, reports that inspection of a CCD spectrum (range 560–860 nm) of SN 1999br, obtained on Apr. 15 by D. Stern (Berkeley) and M. Reuland (Leiden) with the 3m Shane reflector at Lick Observatory, reveals that the object is indeed a supernova, possibly of type II. However, the spectrum is peculiar, with emission lines at observed wavelengths near 580, 720, and 790 nm whose strength is comparable to that of the line near 660 nm (presumably Hα). If the supernova is indeed associated with NGC 4900, it is very subluminous, regardless of its type.

W. Li, University of California at Berkeley, reports that he has searched the KAIT database for observations of NGC 4900 (cf. IAUC 7141), and found no hint of SN 1999br on a poor frame (limiting mag 17) taken on Apr. 4.4 UT. Also, a frame taken on Mar. 27.4 shows no image of the supernova to limiting mag ∼ 18.5.

SUPERNOVA 1999bs IN UGC 11093 Garnavich et al. also report that a spectrum of SN 1999bs (cf. IAUC

7141) was obtained by Calkins on Apr. 15.5 UT. The spectrum consists of a very blue continuum with weak absorption features at Hβ and Hγ, and possibly a broad Hα emission. SN 1999bs, like SN 1999br, is probably a type-II event caught at a very early phase. CCD images obtained by L. Macri with the Fred L. Whipple Observatory 1.2-m telescope, at the same time as the spectra, give V = 16.8.

SUPERNOVA 1999bg IN IC 758 CCD magnitude estimates, unfiltered unless otherwise noted: Mar.

31.593 UT, 15.6 (S. Yoshida and K. Kadota, Ageo City, Japan); Apr. 4.829, R = 15.5 (D. Hanžl, Brno, Czech Republic); 12.845, R = 15.45 (Hanžl); 12.880, V = 15.83 (Hanžl).

1999 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green